OpenStudy (hitaro9):

Find the smallest positive integer k so that 1111x+693y = 10004+k has an integer solution

ganeshie8 (ganeshie8):

start by factoring out the greatest common factor from left hand side

OpenStudy (hitaro9):

Alright 1 minute

ganeshie8 (ganeshie8):

Nope it shouldn't take that long because \(1111 = 1100+11 = 11(100+1)\)

ganeshie8 (ganeshie8):

(jk) take your time :P

OpenStudy (hitaro9):

Oh well we're supposed to show any gcd we find, can't plug in to a calculator

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{}\hspace{.33em}\\~\\ 1111x+693y = 10004+6 \hspace{.33em}\\~\\ (11\times 101)x+(3^2\times 7\times 11)y = (2^2\times 41\times 61) +k \hspace{.33em}\\~\\ \normalsize \text{note that 11 is a common factor between x and y variables ,so}\hspace{.33em}\\~\\ \left((11\times 101)x+(3^2\times 7\times 11)y \right) \pmod {11} \hspace{.33em}\\~\\ = \left((2^2\times 41\times 61)+k\right) \pmod {11} \hspace{.33em}\\~\\ 0+0=(4\times -3\times -5 +k) \pmod {11} \hspace{.33em}\\~\\ 0=(5 +k) \pmod {11} \hspace{.33em}\\~\\ k=-5 \pmod {11} \hspace{.33em}\\~\\ k=6 \pmod {11} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (hitaro9):

Okay I see, each step makes sense as I look through it, thank you.

OpenStudy (mathmath333):

i hope ganeshi's method was same

ganeshie8 (ganeshie8):

Exact same :) In general we have this : \[ax+by = c\] is solvable if and only if \(\gcd(a,b) | c\)

OpenStudy (mathmath333):

what is the name for this corollarry/theorem \(\uparrow \)

ganeshie8 (ganeshie8):

For example 6x + 4y = 10 is solvable in integers because gcd(6,4) = 2 and 2 | 10

ganeshie8 (ganeshie8):

similarly 6x + 4y = 7 is not solvable in integers because gcd(6,4) = 2 and 2\(\nmid\)7

ganeshie8 (ganeshie8):

there is no name as such...

OpenStudy (mathmath333):

ohh

OpenStudy (mathmath333):

i hope someone publishes a paper in int congerence to give a name to it.

ganeshie8 (ganeshie8):

we can call it "solveability of linear congruence theorem" :)

ganeshie8 (ganeshie8):

or "solveability of linear diophantine equation theorem" :)

ganeshie8 (ganeshie8):

as you can see thats too long a name for a simple fact : if \(a*b = c\) it is easy to see conclude that \(a|c\) and \(b|c\)

ganeshie8 (ganeshie8):

\(2*3 = 6\) so we have \(2|6\) and \(3|6\) no need to justify it by appealing to a specific theorem name haha!